Companies with excess cash often employ share repurchase plans in place of or along with cash dividends



9. Stock repurchases 

Companies with excess cash often employ share repurchase plans in place of or along with cash dividends. Share repurchase plans can help investors liquidate their holdings by selling their stock to the issuing company and earning from capital gains. 


Consider the case of Sixty-second Avenue Company. 

Sixty-second Avenue Company has forecasted a net income of $4,200,000 for this year. Its common stock currently trades at $21 per share, and the company currently has 720,000 shares of common stock outstanding. It has sufficient funds available to pay a cash dividend, but many of its investors don't like the additional tax liability to which the dividend income subjects them.


As a result, Sixty-second Avenue's management is considering making a share repurchase transaction in whichit would buy back 75,000 shares of its outstanding shares in the open market by paying the current market share price. Assume that the repurchase transaction will have no effect on either the company's net income or its price-to-earnings (P/E) ratio. What is Sixty-second Avenue's expected stock price after the stock repurchase transaction? 

$29.30 per share 

$21.10 per share 

$23.44 per share 

$28.13 per share 


Which of these factors are considered an advantage of a stock repurchase? Check all that apply. 

Repurchases can be used to produce large-scale changes in capital structure. 

Stockholders who sell their stock back to the implications of the repurchase. 

A repurchase can remove a large block of stock that they were not made fully aware of all that is overhanging the market and keeping the per-share price depressed.
Answer #1

Q-1

Here the price per share is $21 and the company currently have 720000 shares outstanding. Net income is 4200000

EPS of the company will be 4200000/720000 = 5.83

P/E ratio = 21/5.83 = 3.6

After the company buybacks the share, the no of share outstanding will decrease hence it will become 720000 - 75000 = 645000

New EPS will be 4200000/645000 = 6.511

The profit of the company remains the same and also P/E ratio should not change, hence the new price will be,

6.511*3.6 = $23.44

Hence the answer is $23.44

Q-2 The advantage of the stock repurchase is it can produce large scale change in the capital structure of the company as the equity of the company gets reduced the debt/equity ratio changes hence the capital structure changes.